OpenStudy (anonymous):

divide & simplify x²-64 x+8 ----- ÷ ----- x x+7

OpenStudy (anonymous):

x²-64 x+7 ----- ÷ ----- x x+8 (x-8)(x+8) x+7 =---------- ÷ ----- x x+8 x-8 x+7 =----- ÷ ----- x 1

OpenStudy (anonymous):

erm, my bad, but my division is actually suppose to be multiplication

OpenStudy (anonymous):

ok so what is the answer

OpenStudy (anonymous):

well, the problem is practically done, but the reduced form is (x-8) (x+7) =----- * -------- x 1

OpenStudy (anonymous):

it can't be reduced past this point

OpenStudy (anonymous):

ok so then then textbook answer would be (x-8)(x+7) --------- x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x²-64 x+8 ----- ÷ ----- x x+7 (x-8)(x+8) (x+7) =---------- ----- x x+8 (x-8 )( x+7) = -------- --- (x+8) canxel x